OpenStudy (anonymous):

Solve Ths One :) Solve the equation 2(3x + 7) = -22.

OpenStudy (anonymous):

3x+7=-22/2 3x+7=-11 3x=-18 x=-6

OpenStudy (anonymous):

A. x = 6 B. x = -6 C. x = 2 D. x = -2

OpenStudy (anonymous):

Solve this one by yourself